OpenStudy (kayluh123):

If you are in a hot-air balloon at a height of 100 meters, you can expect to see ____ km. Round your answer to the nearest kilometer and enter only the number.

OpenStudy (anonymous):

if earth's radius is r meters, and distance to horizon is d, r^2 + d^2 = (r+100)^2 plug in r and solve for d.

OpenStudy (anonymous):

use conversion technique called dimensional analysis. 100m * 1km/m = 100km

OpenStudy (anonymous):

but that is just the distance from the balloon to the horizon. The actual distance seen along the earth is the part of the circumference subtended by the angle t, where cos(t) = r/(r+h) the distance measured along the surface of the earth is then d = r*arccos(r/(r+h))